A ballistic pendulum is a device that is used to determine the muzzle velocity of the bullet shot by a hunting rifle, for example. As you know, the velocity of the bullet fired by a rifle dictates the distance it can travel. Those who reload their own bullets may find this device helpful. When a bullet is fired broadside into a suspended 4x4 chunk of wood, the wood swings up from rest by an angle of 60 degrees. Together, the bullet and chuck of wood has a mass of 2 kg. For supporting wires in length of 50 cm, a) show that as a result, the block swings up vertically a distance of 0.25 m, and b) correspondingly the woods change of potential energy is 4.9 Joules. For a bullet of mass 0.01 kg, what is the muzzle velocity of the bullet?
Answer #1

1) So vertical swung by bullet and wood block is

h= lcos\theta =(50 cm) cos60= 25 cm= 0.25 m2) So Potential Energy is

P = mgh

So,

P = 2× 9.8× 0.25 = 4.9 J

3)So from energy conservation P.E = K.E

So,

4.9 =\frac{1}{2} mv^2

So, m= 0.01kg

We get,

4.9 =\frac{1}{2} (0.01)v^2

v = 31.3 m/s